# Parameter-Efficient Fine-Tuning of State Space Models

### v1

@@ C.4 Optimal Application of LoRA in SSM-based Models
$=\sum_{m=1}^{n}\operatorname{diag}\left({\bm{P}}^{\top}\overline{{\bm{A}}}{\bm{P}}\right)^{n-m}\left(({\bm{P}}^{\top}{\bm{C}}^{\top})\otimes({\bm{P}}^{\top}\overline{{\bm{B}}})\right)x_{m}$ (45)
Therefore, to make the updated model equivalent to the target model, we need to update ${\bm{P}}^{\top}\overline{{\bm{A}}}{\bm{P}}$ to align with $\begin{bmatrix}\overline{{\bm{A}}}_{\star}&{\bm{O}}\\ {\bm{O}}&{\bm{O}}\\ \end{bmatrix}$, and $({\bm{P}}^{\top}{\bm{C}}^{\top})\otimes({\bm{P}}^{\top}\overline{{\bm{B}}})$ to align with $\begin{bmatrix}{\bm{C}}_{\star}^{\top}\\ {\bm{0}}\end{bmatrix}\otimes\begin{bmatrix}\overline{{\bm{B}}}_{\star}\\ {\bm{0}}\end{bmatrix}$. If they are already matching or partially matched for certain entries, no updates are required for those entries; only the unmatched entries need to be updated. Then, the required trainable parameters for this permutation matrix ${\bm{P}}$ are:
$\left\lVert\left[{\bm{P}}^{\top}(\operatorname{diag}(\overline{{\bm{A}}})\otimes\overline{{\bm{B}}}\otimes{\bm{C}}^{\top})\right]_{(H_{\star}+1):H}\right\rVert_{0}+\left\lVert\left[{\bm{P}}^{\top}\overline{{\bm{A}}}{\bm{P}}\right]_{1:H_{\star},1:H_{\star}}-\overline{{\bm{A}}}_{\star}\right\rVert_{0}+\left\lVert\left[{\bm{P}}^{\top}(\overline{{\bm{B}}}\otimes{\bm{C}}^{\top})\right]_{1:H_{\star}}-\overline{{\bm{B}}}_{\star}\otimes{\bm{C}}_{\star}^{\top}\right\rVert_{0}.$
@@
This lemma highlights the significance of identifying essential hidden state dimensions. The term $\left\lVert\left[{\bm{P}}^{\top}(\operatorname{diag}(\overline{{\bm{A}}})\otimes\overline{{\bm{B}}}\otimes{\bm{C}}^{\top})\right]_{(H_{\star}+1):H}\right\rVert_{0}$ underscores the importance of excluding redundant dimensions. This can be achieved by either directly removing these dimensions from the state matrix $\overline{{\bm{A}}}$, or by updating $\overline{{\bm{B}}}$ or ${\bm{C}}$ to ensure that only the selected hidden state dimensions are utilized during the input transition or output mapping phases. Once redundant dimensions are filtered out, tuning only the essential dimensions is sufficient to align the updated model with the target model.
Furthermore, based on the lemma, the roles of the input transition vector $\overline{{\bm{B}}}$ and ${\bm{C}}^{\top}$ are nearly identical, as they consistently appear together as the combined term $\overline{{\bm{B}}}\otimes{\bm{C}}^{\top}$, which is also discussed in Gupta et al. (2022). Consequently, one could opt to tune either $\overline{{\bm{B}}}$ or ${\bm{C}}$ exclusively or alternatively, split the indices into two groups, tuning $\overline{{\bm{B}}}$ for the first group and ${\bm{C}}$ for the second. Both vectors indicate how information from different hidden state dimensions is integrated, whereas $\overline{{\bm{A}}}$ plays a distinct role, determining how the hidden states are stored.
In practice, instead of directly using the discretized parameters $\overline{{\bm{A}}},\overline{{\bm{B}}},{\bm{C}}$, S4 is implemented using the continuous parameters ${\bm{A}},{\bm{B}},{\bm{C}}$ with step size $\Delta$. To provide further practical guidance on parameter tuning, the following two lemmas analyze the parameter efficiency of continuous parameters under different discretization methods: Two exemplary methods of discretization are bilinear and zero-order hold (ZOH):
@@
Consider the parameters before discretization, i.e., ${\bm{A}},{\bm{B}},{\bm{C}}$, and they are discretized via bilinear discretization. To achieve functional equivalence between the updated model and the target model, i.e., $\hat{f}\equiv f^{\star}$, it is sufficient to tune the following number of parameters:
$\min_{{\bm{P}}\in{\mathcal{P}}^{H}}\overbrace{\left\lVert\left[\Delta{\bm{P}}^{\top}(\operatorname{diag}({\bm{I}}+\Delta/2{\bm{A}})\otimes{\bm{B}}\otimes{\bm{C}}^{\top})\right]_{(H_{\star}+1):H}\right\rVert_{0}}^{\text{eliminating redundant dimensions}}+\overbrace{\underbrace{\left\lVert\left[{\bm{P}}^{\top}{\bm{A}}{\bm{P}}\right]_{1:H_{\star},1:H_{\star}}-{\bm{A}}_{\star}\right\rVert_{0}}_{\text{aligning the state matrix}}+\underbrace{\left\lVert\left[{\bm{P}}^{\top}({\bm{B}}\otimes{\bm{C}}^{\top})\right]_{1:H_{\star}}-{\bm{B}}_{\star}\otimes{\bm{C}}_{\star}^{\top}\right\rVert_{0}}_{\text{aligning input-output interactions}}}^{\text{aligning used dimensions with target model}}.$
@@
Consider the parameters before discretization, i.e., ${\bm{A}},{\bm{B}},{\bm{C}}$, and they are discretized via ZOH discretization. To achieve functional equivalence between the updated model and the target model, i.e., $\hat{f}\equiv f^{\star}$, it is sufficient to tune the following number of parameters:
$\min_{{\bm{P}}\in{\mathcal{P}}^{H}}\overbrace{\left\lVert\left[\Delta{\bm{P}}^{\top}(\operatorname{diag}(\exp(\Delta{\bm{A}})-{\bm{I}})\otimes{\bm{B}}\otimes{\bm{C}}^{\top})\right]_{(H_{\star}+1):H}\right\rVert_{0}}^{\text{eliminating redundant dimensions}}+\overbrace{\underbrace{\left\lVert\left[{\bm{P}}^{\top}{\bm{A}}{\bm{P}}\right]_{1:H_{\star},1:H_{\star}}-{\bm{A}}_{\star}\right\rVert_{0}}_{\text{aligning the state matrix}}+\underbrace{\left\lVert\left[{\bm{P}}^{\top}({\bm{B}}\otimes{\bm{C}}^{\top})\right]_{1:H_{\star}}-{\bm{B}}_{\star}\otimes{\bm{C}}_{\star}^{\top}\right\rVert_{0}}_{\text{aligning input-output interactions}}}^{\text{aligning used dimensions with target model}}.$
@@
The insights provided by Lemma 7 and Lemma 8 are the same as those provided by Lemma 3. The analysis here supports the second step of SDLoRA presented in Sec. 5.
@@ D.2 Extension to Deep S4 Models
Our previous analysis focused on single-channel S4 models. We now expand our investigation to more complex scenarios involving deep S4 models for both target and frozen architectures, incorporating $D$ channels and varying layer depths. In this section, we consider two PEFT methods: (i) Selective Dimension Tuning (SDT) and (ii) SDLoRA. The key distinction between SDT and SDLoRA lies in their treatment of linear projection matrices. SDT exclusively updates the columns of weight matrices corresponding to the updatable channels identified through Alg. 1. In contrast, SDLoRA employs LoRA to modify these matrices. It is worth noting that the linear projection matrix updates in SDT are inherently low-rank, making it a specialized case of SDLoRA. Our analysis starts with SDT, and it automatically applies to SDLoRA.
In this analysis, we assume that each input token $x_{t}$ belongs to ${\mathcal{X}}$, a bounded subset of ${\mathbb{R}}^{D}$, and that the length of the input sequence is finite. Let the frozen model have $L$ layers, and the target model have $L^{\star}$ layers, where $L\geq L^{\star}$. Similar to the technique used in Zeng & Lee (2024) and Giannou et al. (2023). The basic idea of updating the frozen model to match the functionality of the target model is to utilize every $\lceil L/L^{\star}\rceil$ layers of the frozen model to approximate every layer of the target model. We start introducing this proof idea from the simplest case where $L^{\star}=1,L=D$. In this scenario, we can simply choose one different channel to tune and maintain all other channels at zero at every layer. The outputs from the various channels of the deep S4 layers are then combined through a residual connection. This proof idea inspires us to perform channel selection and make use of the residual connections, which is the first and third step of SDLoRA presented in Sec. 5. Building on this idea, we present the following results for when the target model has only $L^{\star}=1$ layer, and $L=D=2$.
@@
Consider a $D$-dimensional input sequence. Assume that the linear layers in the model have linear activation functions. Using SDT, any deep S4 model with $H$ hidden states per channel and $L$ layers can be updated to accurately present any target one-layer deep S4 model without residual connections, having a reduced hidden state dimension $H^{\star}<H$. Then this can be achieved by selectively fine-tuning at most $\lceil D/L\rceil$ channels, $H^{\star}$ hidden states, and residual connections at each layer, while additionally fully fine-tuning the linear projection matrix of the last layer only.
@@
$\begin{split}{\bm{z}}_{l,t}&=\widehat{{\bm{W}}}_{l}\cdot\widehat{\operatorname{S4}}_{l,t}({\bm{z}}_{l-1,1},\ldots,{\bm{z}}_{l-1,t})+\widehat{{\bm{\beta}}}_{l}+\widehat{{\bm{u}}}_{l}\otimes{\bm{z}}_{l-1,t}\\ &=\widehat{{\bm{W}}}_{l}\cdot\begin{bmatrix}\widehat{\operatorname{S4}}_{l,t}^{(1)}(z_{l-1,1}^{(1)},\ldots,z_{l-1,t}^{(1)})\\ \vdots\\ \widehat{\operatorname{S4}}_{l,t}^{(D)}(z_{l-1,1}^{(D)},\ldots,z_{l-1,t}^{(D)})\\ \end{bmatrix}+\widehat{{\bm{\beta}}}_{l}+\widehat{{\bm{u}}}_{l}\otimes{\bm{z}}_{l-1,t},\end{split}$ (50)
@@
  • (Residual and Bias Tuning) Regardless of other selections, SDLoRA consistently tunes the coefficients of residual connections and biases in linear projections, as these components contain a negligible number of parameters. In this scenario, we let $\widehat{{\bm{\beta}}}_{l}={\bm{0}},\leavevmode\nobreak\ \widehat{{\bm{u}}}_{l}=\begin{bmatrix}\smash{\underbrace{\begin{matrix}1&\cdots&1\end{matrix}}_{l-1\text{ elements}}}&0&\smash{\underbrace{\begin{matrix}1&\cdots&1\end{matrix}}_{D-l\text{ elements}}}\end{bmatrix}^{\top}.$
@@
Theorem 10 (Expressive Power of SDLoRA on Deep S4 Models).
Consider a $D$-dimensional input sequence. Assume that the linear layers in the model have linear activation functions. Using SDT, any deep S4 model with $H$ hidden states per channel and $L$ layers can be updated to accurately present any target deep S4 model without residual connections, having a reduced hidden state dimension $H^{\star}<H$, and fewer layers $L^{\star}<L$. This can be achieved by selectively fine-tuning at most $\lceil DL^{\star}/L\rceil$ channels, $H^{\star}$ hidden states, and residual connections at each layer.
@@
For reader’s convience, we restate the following statement presented in the main body again here.
Theorem 0 (Expressive Power of SDLoRA on Deep S4 Models).
Consider a $D$-dimensional input sequence. Assume that the linear layers in the model have linear activation functions. Using SDLoRA, any deep S4 model with $H$ hidden states per channel and $L$ layers can be updated to accurately present any target deep S4 model without residual connections, having a reduced hidden state dimension $H^{\star}<H$, and fewer layers $L^{\star}<L$. This can be achieved by selectively fine-tuning at most $\lceil DL^{\star}/L\rceil$ channels, $H^{\star}$ hidden states on SSM modules, applying rank-$\lceil\frac{L}{L^{\star}}\rceil$ updates on linear projection matrices and updating residual connections and biases at each layer, while additionally fully fine-tuning the linear projection matrix of the last layer only.

### v3

@@ C.4 Optimal Application of LoRA in SSM-based Models
$=\sum_{m=1}^{n}\operatorname{diag}\left({\bm{P}}^{\top}\overline{{\bm{A}}}{\bm{P}}\right)^{n-m}\left(({\bm{P}}^{\top}{\bm{C}}^{\top})\odot({\bm{P}}^{\top}\overline{{\bm{B}}})\right)x_{m}.$ (46)
Due to the convolution structure of $\overline{{\bm{A}}}$, two models are functionally equivalent if and only if ${\bm{P}}^{\top}\overline{{\bm{A}}}{\bm{P}}$ aligns with $\begin{bmatrix}\overline{{\bm{A}}}_{\star}&{\bm{O}}\\ {\bm{O}}&{\bm{O}}\\ \end{bmatrix}$, and $({\bm{P}}^{\top}{\bm{C}}^{\top})\odot({\bm{P}}^{\top}\overline{{\bm{B}}})$ align with $\begin{bmatrix}{\bm{C}}_{\star}^{\top}\\ {\bm{0}}\end{bmatrix}\odot\begin{bmatrix}\overline{{\bm{B}}}_{\star}\\ {\bm{0}}\end{bmatrix}$ for some $P\in{\mathcal{P}}^{H}$. If they are already matching or partially matched for certain entries, no updates are required for those entries; only the unmatched entries need to be updated. Then, the required trainable parameters for this permutation matrix ${\bm{P}}$ are:
$\left\lVert\left[{\bm{P}}^{\top}(\operatorname{diag}(\overline{{\bm{A}}})\odot\overline{{\bm{B}}}\odot{\bm{C}}^{\top})\right]_{(H_{\star}+1):H}\right\rVert_{0}+\left\lVert\left[{\bm{P}}^{\top}\overline{{\bm{A}}}{\bm{P}}\right]_{1:H_{\star},1:H_{\star}}-\overline{{\bm{A}}}_{\star}\right\rVert_{0}+\left\lVert\left[{\bm{P}}^{\top}(\overline{{\bm{B}}}\odot{\bm{C}}^{\top})\right]_{1:H_{\star}}-\overline{{\bm{B}}}_{\star}\odot{\bm{C}}_{\star}^{\top}\right\rVert_{0}.$
@@
This lemma highlights the significance of identifying essential hidden state dimensions. The term $\left\lVert\left[{\bm{P}}^{\top}(\operatorname{diag}(\overline{{\bm{A}}})\odot\overline{{\bm{B}}}\odot{\bm{C}}^{\top})\right]_{(H_{\star}+1):H}\right\rVert_{0}$ underscores the importance of excluding redundant dimensions. This can be achieved by either directly removing these dimensions from the state matrix $\overline{{\bm{A}}}$, or by updating $\overline{{\bm{B}}}$ or ${\bm{C}}$ to ensure that only the selected hidden state dimensions are utilized during the input transition or output mapping phases. Once redundant dimensions are filtered out, tuning only the essential dimensions is sufficient to align the updated model with the target model.
Furthermore, based on the lemma, the roles of the input transition vector $\overline{{\bm{B}}}$ and ${\bm{C}}^{\top}$ are nearly identical, as they consistently appear together as the combined term $\overline{{\bm{B}}}\odot{\bm{C}}^{\top}$, which is also discussed in Gupta et al. (2022). Consequently, one could opt to tune either $\overline{{\bm{B}}}$ or ${\bm{C}}$ exclusively or alternatively, split the indices into two groups, tuning $\overline{{\bm{B}}}$ for the first group and ${\bm{C}}$ for the second. Both vectors indicate how information from different hidden state dimensions is integrated, whereas $\overline{{\bm{A}}}$ plays a distinct role, determining how the hidden states are stored.
In practice, instead of directly using the discretized parameters $\overline{{\bm{A}}},\overline{{\bm{B}}},{\bm{C}}$, S4 is implemented using the continuous parameters ${\bm{A}},{\bm{B}},{\bm{C}}$ with step size $\Delta$. To provide further practical guidance on parameter tuning, the following two lemmas analyze the parameter efficiency of continuous parameters under different discretization methods. Two exemplary methods of discretization are bilinear and zero-order hold (ZOH):
@@
Consider the parameters before discretization, i.e., ${\bm{A}},{\bm{B}},{\bm{C}}$, which are subsequently discretized using bilinear discretization. To achieve functional equivalence between the updated model and the target model, i.e., $\hat{f}\equiv f^{\star}$, it is sufficient to tune the following number of parameters:
$\min_{{\bm{P}}\in{\mathcal{P}}^{H}}\overbrace{\left\lVert\left[\Delta{\bm{P}}^{\top}(\operatorname{diag}({\bm{I}}+\Delta/2{\bm{A}})\odot{\bm{B}}\odot{\bm{C}}^{\top})\right]_{(H_{\star}+1):H}\right\rVert_{0}}^{\text{eliminating redundant dimensions}}+\overbrace{\underbrace{\left\lVert\left[{\bm{P}}^{\top}{\bm{A}}{\bm{P}}\right]_{1:H_{\star},1:H_{\star}}-{\bm{A}}_{\star}\right\rVert_{0}}_{\text{aligning the state matrix}}+\underbrace{\left\lVert\left[{\bm{P}}^{\top}({\bm{B}}\odot{\bm{C}}^{\top})\right]_{1:H_{\star}}-{\bm{B}}_{\star}\odot{\bm{C}}_{\star}^{\top}\right\rVert_{0}}_{\text{aligning input-output interactions}}}^{\text{aligning used dimensions with target model}}.$
@@
Consider the parameters before discretization, i.e., ${\bm{A}},{\bm{B}},{\bm{C}}$, which are subsequently discretized using ZOH discretization. To achieve functional equivalence between the updated model and the target model, i.e., $\hat{f}\equiv f^{\star}$, it is sufficient to tune the following number of parameters:
$\min_{{\bm{P}}\in{\mathcal{P}}^{H}}\overbrace{\left\lVert\left[\Delta{\bm{P}}^{\top}(\operatorname{diag}(\exp(\Delta{\bm{A}})-{\bm{I}})\odot{\bm{B}}\odot{\bm{C}}^{\top})\right]_{(H_{\star}+1):H}\right\rVert_{0}}^{\text{eliminating redundant dimensions}}+\overbrace{\underbrace{\left\lVert\left[{\bm{P}}^{\top}{\bm{A}}{\bm{P}}\right]_{1:H_{\star},1:H_{\star}}-{\bm{A}}_{\star}\right\rVert_{0}}_{\text{aligning the state matrix}}+\underbrace{\left\lVert\left[{\bm{P}}^{\top}({\bm{B}}\odot{\bm{C}}^{\top})\right]_{1:H_{\star}}-{\bm{B}}_{\star}\odot{\bm{C}}_{\star}^{\top}\right\rVert_{0}}_{\text{aligning input-output interactions}}}^{\text{aligning used dimensions with target model}}.$
@@
The insights provided by Lemma 3 and Lemma 4 are the same as those provided by Lemma 2. The analysis here supports the second step of SDT-P presented in Sec. 5.
@@ D.2 Extension to Deep S4 Models
Our previous analysis focused on single-channel S4 models. We now expand our investigation to more complex scenarios involving deep S4 models for both target and frozen architectures, incorporating $D$ channels and varying layer depths. In this section, in addition to SDT-P, we introduce SDT+. The key distinction between SDT+ and SDT-P lies in their treatment of linear projection matrices. SDT-P only operate on SSM modules and additionally requires applying LoRA to modify the linear projection matrices. In contrast, SDT+ applies SDT-P on SSM modules while also updates the columns of weight matrices corresponding to the updatable channels identified through Alg. 2. It is worth noting that the linear projection matrix updates in SDT+ are inherently low-rank, making it a specialized case of SDT-P combined with LoRA. Our analysis starts with SDT+, and it automatically applies to SDT-P combined with LoRA.
In this analysis, we assume that each input token $x_{t}$ belongs to ${\mathcal{X}}$, a bounded subset of ${\mathbb{R}}^{D}$, and that the length of the input sequence is finite. Let the frozen model have $L$ layers, and the target model have $L^{\star}$ layers, where $L\geq L^{\star}$. Similar to the technique used in Zeng & Lee (2024) and Giannou et al. (2023). The basic idea of updating the frozen model to match the functionality of the target model is to utilize every $\lceil L/L^{\star}\rceil$ layers of the frozen model to approximate every layer of the target model. We start introducing this proof idea from the simplest case where $L^{\star}=1,L=D$. In this scenario, we can simply choose one different channel to tune and maintain all other channels at zero at every layer. The outputs from the various channels of the deep S4 layers are then combined through a residual connection. This proof idea inspires us to perform channel selection and make use of the residual connections, which is the first and third step of SDT-P presented in Sec. 5. Building on this idea, we present the following results for when the target model has only $L^{\star}=1$ layer, and $L=D=2$.
@@
Consider a $D$-dimensional input sequence. Assume that the linear layers in the model have linear activation functions. Using SDT+, any deep S4 model with $H$ hidden states per channel and $L$ layers can be updated to accurately present any target one-layer deep S4 model without residual connections, having a reduced hidden state dimension $H^{\star}<H$. Then this can be achieved by selectively fine-tuning at most $\lceil D/L\rceil$ channels, $H^{\star}$ hidden states, and residual connections at each layer, while additionally fully fine-tuning the linear projection matrix of the last layer only.
@@
$\begin{split}{\bm{z}}_{l,t}&=\widehat{{\bm{W}}}_{l}\cdot\widehat{\operatorname{S4}}_{l,t}({\bm{z}}_{l-1,1},\ldots,{\bm{z}}_{l-1,t})+\widehat{\bm{\beta}}_{l}+\widehat{{\bm{u}}}_{l}\odot{\bm{z}}_{l-1,t}\\ &=\widehat{{\bm{W}}}_{l}\cdot\begin{bmatrix}\widehat{\operatorname{S4}}_{l,t}^{(1)}(z_{l-1,1}^{(1)},\ldots,z_{l-1,t}^{(1)})\\ \vdots\\ \widehat{\operatorname{S4}}_{l,t}^{(D)}(z_{l-1,1}^{(D)},\ldots,z_{l-1,t}^{(D)})\\ \end{bmatrix}+\widehat{\bm{\beta}}_{l}+\widehat{{\bm{u}}}_{l}\odot{\bm{z}}_{l-1,t},\end{split}$ (51)
@@
  • (Residual and Bias Tuning) Regardless of other selections, SDT consistently tunes the coefficients of residual connections and biases in linear projections, as these components contain a negligible number of parameters. In this scenario, we let $\widehat{\bm{\beta}}_{l}={\bm{0}},\leavevmode\nobreak\ \widehat{{\bm{u}}}_{l}=\begin{bmatrix}\smash{\underbrace{\begin{matrix}1&\cdots&1\end{matrix}}_{l-1\text{ elements}}}&0&\smash{\underbrace{\begin{matrix}1&\cdots&1\end{matrix}}_{D-l\text{ elements}}}\end{bmatrix}^{\top}.$
@@
Theorem 2 (Expressive Power of SDT+ on Deep S4 Models).
Consider a $D$-dimensional input sequence. Assume that the linear layers in the model have linear activation functions. Using SDT+, any deep S4 model with $H$ hidden states per channel and $L$ layers can be updated to accurately present any target deep S4 model without residual connections, having a reduced hidden state dimension $H^{\star}<H$, and fewer layers $L^{\star}<L$. This can be achieved by selectively fine-tuning at most $\lceil DL^{\star}/L\rceil$ channels, $H^{\star}$ hidden states, and residual connections at each layer.
@@
Theorem 2 leads to the following result, which represents the deep S4 model case of Theorem 1.
Theorem 3 (Expressive Power of SDT-P on Deep S4 Models).
Consider a $D$-dimensional input sequence. Assume that the linear layers in the model have linear activation functions. Using SDT-P, any deep S4 model with $H$ hidden states per channel and $L$ layers can be updated to accurately present any target deep S4 model without residual connections, having a reduced hidden state dimension $H^{\star}<H$, and fewer layers $L^{\star}<L$. This can be achieved by selectively fine-tuning at most $\lceil DL^{\star}/L\rceil$ channels, $H^{\star}$ hidden states on SSM modules, applying rank-$\lceil\frac{L}{L^{\star}}\rceil$ updates on linear projection matrices and updating residual connections and biases at each layer, while additionally fully fine-tuning the linear projection matrix of the last layer only.